OpenStudy (anonymous):

If (x+y)^(1/2) = (x-y)^(1/2) which of the following must be true? y=0 x+y=1 x-y=1 x^2-y^2 =1 none of the above

OpenStudy (amistre64):

sqrt(x+y) = sqrt(x-y)? y=0 is true.